You have the following data: FCF0 = $10 million; FCF1 = $15 million; FCF2 = $20 million; FCF3 = $25 million; free cash flow grows at a rate of 5% for year 4 and beyond. The weighted average cost of capital is 15%. Assume they have 40 million in debt and 10 million shares outstanding. Find the price per share.
1
Expert's answer
2012-12-19T11:00:37-0500
Present value of FCF0-3 is PV=FCF0+FCF1/(1+0.15)+FCF2/(1+0.15)^2+FCF3/(1+0.15)^3=54.6
Then using formula for permanent growth we get value of all in flows which are expected to appear after FCF3 at moment of time t=3:
PV3=FCF3/(0.15-0.05)=250. Than we discount it to the current time PV=PV3/(1+0.15)^3=164.4.
Total value of all shares: TV=54.6+164.4-40=179
PPS=179/10=17.9 $ per share.